OpenStudy (anonymous):

Help me !!!! Find the volume of a prism with triangular bases where the perimeter of the base = 18", apothem = x", and the prism height = 12". USe V =Bh and A = 1/2 ap for the area of the base.

OpenStudy (anonymous):

can you draw it?

OpenStudy (ranga):

A = 1/2 * a * p ; where A is the area of the case; a is apothem and p is perimeter p = 18, a = x. Substitute in A Area A = 1/2*18 * x = 9x Volume of prism = Area of base * height = 9x * 12 = 108x